OpenStudy (anonymous):

What is the slope of the line 5x + 1y = 6?

OpenStudy (anonymous):

(1,1) ?

OpenStudy (anonymous):

Solve the equation for y: y=-5x+6 This function follows the form y=mx+b (ie the equation for a line, where m is the slope and b is the y-intercept) Can you find the slope now?

OpenStudy (anonymous):

1 !

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

fantastic haha

OpenStudy (anonymous):

The slope of a line will always be the coefficient of x. What is the coefficient of x in the equation: y=-5x+6 If you can identify the coefficient of x in that equation you can identify the slope of the line

OpenStudy (anonymous):

-5......

OpenStudy (anonymous):

That is correct. If you are having trouble with the concept, use the idea that the slope is (y2-y1)/(x2-x1), colloquially known as "rise over run." Write out a table of values for the function and plot them on a graph, ie: for x=1, y=-5(1)+6=1 for x=2, y=-5(2)+6=-4 for x=4, y=-5(4)+6=-14 Take any two points from your graph and plug them into the formula "rise over run" I will take x=1 and x=4 (y2-y1)/(x2-x1)=(-14-1)/(4-1)=(-15)/3=-5 So the slope is -5

OpenStudy (raden):

another way : ax+by=c, the slope is m = -a/b so, if given 5x + 1y = 6 then m = -5/1 = -5